# Modeling the Multivariate Relationship with Contextualized Representations for Effective Human-Object Interaction Detection

### v1

@@ Abstract
Human-Object Interaction (HOI) detection aims to simultaneously localize human-object pairs and recognize their interactions. While recent two-stage approaches have made significant progress, they still face challenges due to incomplete context modeling. In this work, we introduce a Contextualized Representation Learning Network that integrates both affordance-guided reasoning and contextual prompts with visual cues to better capture complex interactions. We enhance the conventional HOI detection framework by expanding it beyond simple human-object pairs to include multivariate relationships involving auxiliary entities like tools. Specifically, we explicitly model the functional role (affordance) of these auxiliary objects through triplet structures <<human, tool, object>>. This enables our model to identify tool-dependent interactions such as “filling”. Furthermore, the learnable prompt is enriched with instance categories and subsequently integrated with contextual visual features using an attention mechanism. This process aligns language with image content at both global and regional levels. These contextualized representations equip the model with enriched relational cues for more reliable reasoning over complex, context-dependent interactions. Our proposed method demonstrates superior performance on both the HICO-Det and V-COCO datasets in most scenarios. Codes will be released upon acceptance.
Introduction
@@
Human-Object Interaction (HOI) detection is a fundamental and challenging task in visual recognition that aims to model complex semantic relationships between humans and surrounding objects in a visual scene, ultimately recognizing triplets of the form <<human, action, object>>. It requires a simultaneous understanding of visual features, spatial configurations and contextual semantics to effectively identify all valid interaction patterns within given images. Therefore, other high-level semantic understanding tasks, such as activity recognition (Wang et al. 2025) and video comprehension (Tao et al. 2024) can benefit from HOI.
@@
Figure 1: Contextualized representations in HOI. (a) Tool affordances (e.g. bottle’s pourable) help distinguish complex interactions (e.g., human fill cup) from direct human–object relations (e.g., human hold cup). (b) Contextualized alignment with instance categories (e.g., cup) and corresponding visual features narrow down the potential actions to relevant ones (e.g., fill, hold).
@@
Existing HOI detection approaches can generally be divided into one-stage and two-stage methods. One-stage methods perform object detection and interaction prediction simultaneously, but adapting their decoders for various HOIs often incurs substantial training costs, sometimes requiring hundreds of GPU hours. On the other hand, two-stage methods leverage pre-trained object detectors to identify humans and objects before explicitly establishing relationships between them for interaction prediction. Thanks to its flexibility, the two-stage approach has quickly become the preferred paradigm. However, traditional two-stage HOI approaches have primarily relied on features derived from human-object pairs, such as human pose (Wu et al. 2024) or spatial information (Zhang, Campbell, and Gould 2022). In real-world interactions, many scenarios are not solely defined by these pairs. They may also involve auxiliary tools or objects that play a crucial role in the interaction. For example, as illustrated in Fig. 1(a), when a person fills a cup using a bottle, understanding the bottle’s affordance (i.e. its potential function like being pourable) offers significant insights into deciphering the nature of the interaction.
@@
Meanwhile, recent advancements in pre-trained language models like BERT (Devlin et al. 2019) and GPT (Brown et al. 2020) offer rich semantic insights for understanding HOIs. Prompt Learning, as an efficient method for finetuning language models, has been extened to visual tasks by CoOp (Zhou et al. 2022b). Many studies have as integrated Prompt Learning into the HOI task (Wang et al. 2022; Lei et al. 2024; Lei, Wang, and Tan 2024). However, these approaches rely exclusively on textual data and do not integrate image information from specific samples. As a result, crucial contextual visual cues are left untapped, which are essential for accurately capturing interactions.
@@
It’s worthy noting that both the affordances and the text-related appearance of a specific object act as vital forms of contextualized representation, essential for identifying complex interactions. Despite the importance, their applications in HOI has not been thoroughly explored. This gap in research greatly motivates us to delve deeper into this area. In this work, we propose a new two-stage HOI detection framework, named Contextualized Representation Learning Network (CRLN). To be specific, unlike traditional human-object pairs, multivariate relationships (unary, binary and ternary ones) are modeled to capture tool-mediated interactions through affordance-guided triplets <<human, tool, object>>. In parallel, the contextualized representations of detected instances are injected into the process of prompt learning. Specifically, as illustrated in Fig. 1(b), we extract localized visual features from the corresponding instance regions and refine the prompt representation through a cross-attention mechanism with VLM’s global image embedding. This design allows the prompt to be conditioned on both textual semantics and instance-level visual cues, improving its capacity to capture context-dependent interactions.
@@
Related Work
Human-Object Interaction Detection
@@
Existing Human-Object Interaction (HOI) detection approaches can be divided into one-stage and two-stage methods. One-stage methods (Zou et al. 2021; Kim et al. 2021; Tamura, Ohashi, and Yoshinaga 2021; Xie et al. 2023) simultaneously predict human and object boxes, categories, and interaction classes. Recent advancements in one-stage detectors, particularly those utilizing transformer architectures (Zhang et al. 2021; Kim, Jung, and Cho 2023; Zhang et al. 2024; Gao et al. 2024), have shown promising performance. This end-to-end approach simplifies the inference process of HOI. However, it imposes heavy computational costs during training, which can limit its practicality.
@@
To alleviate this issue, two-stage methods (He et al. 2021; Liu et al. 2022; Wu et al. 2022; Zhang et al. 2022; Zheng, Xu, and Jin 2023; Guo et al. 2024) undergone significant development recently. They typically rely on pre-trained detectors (e.g. DETR) to generate object proposals, focusing more on extracting interaction context from candidate human-object pairs. For example, UPT (Zhang, Campbell, and Gould 2022) first detects all humans and objects (i.e. unary features) and then applies self-attention to unary features and human-object pairs, effectively enhancing the confidence of positive samples. PViC (Zhang et al. 2023) incorporates image features back into the representation of human-object pairs via cross-attention to compensate for missing contextual information. Nonetheless, traditional methods often treat all objects equally, failing to account for the role that tools play in interactions. In this work, we explore the concept of tool affordances to more effectively capture these relational cues.
Vision-Language Models in HOI
@@
To obtain more effective HOI representations, several studies (Liao et al. 2022) explore knowledge transfer from vision-language pre-trained models (e.g. CLIP (Radford et al. 2021)). This approach not only enriches the learned representations but also improves the model’s capability in HOI detection. GEN-VLKT (Liao et al. 2022) utilizes CLIP knowledge for interaction classification and the distillation of visual features. HOICLIP (Ning et al. 2023) uses the features obtained by the VLM visual encoder and proposes a new transfer strategy that uses visual semantic algorithms to represent action. CLIP4HOI(Mao et al. 2023) utilizes CLIP’s vision-language knowledge to enhance human-object interaction detection by decoupling human and object detection and adapting CLIP into a fine-grained classifier for better interaction discrimination.
Figure 2: Overall architecture of our Contextualized Representation Learning Network, consisting of Multivariate Relationship Modeling (MRM) and Contextualized Prompt Learning (CPL). MRM constructs unary, binary and ternary token sets from regional features to model HOIs. CPL builds a category-aware learnable prompt, fused with diverse contextual visual features. Their combined outputs are utilized for interaction prediction. The structure of binary/ternary/contextual decoder is shown in the bottom right.
Prompt Learning in HOI
@@
Prompt learning has become very popular for fine-tuning VLMs on downstream tasks. Context Optimization (CoOp) (Zhou et al. 2022b) encodes prompt context as learnable vectors, achieving strong performance with only a few labeled samples. To better adapt to downstream tasks, Conditional Context Optimization (CoCoOp) (Zhou et al. 2022a) builds upon CoOp by employing a lightweight network to generate input-conditional context tokens for each image. THID (Wang et al. 2022) is the first to propose the use of learnable language prompts for the HOI detection task. CMMP (Lei et al. 2024) introduces decoupled multi-modal prompts for spatial-aware HOI detection, separating visual feature extraction and interaction classification to reduce error propagation. EZ-HOI (Lei, Wang, and Tan 2024) leverages LLM-generated class descriptions to guide prompt learning.
@@
Methodology
Preliminary
Our proposed model builds upon the traditional two-stage methodology for Human-Object Interaction (HOI) detection, which first focuses on detecting instances and subsequently classifies the interactions between them. In the first stage, as shown in the upper-left part of Fig. 2, we employ an off-the-shelf object detector, such as DETR (Carion et al. 2020), to extract all instances, including humans and objects. As a result, a set of instances $\mathcal{Z}=\{z_{i}\}_{i=1}^{n}$ can be constructed with each detection $z_{i}\,=\,(\mathbf{b}_{i},s_{i},c_{i},\mathbf{u}_{i})$, consisting of the box coordinates $\mathbf{b}_{i}\in\mathbb{R}^{4}$, the confidence score $s_{i}\in[0,1]$, the detected instance $c_{i}\in\mathcal{O}$ for the category set $\mathcal{O}$, and the unary instance feature $\mathbf{u}_{i}\in\mathbb{R}^{C}$. $C$ is the embedding dimension, $n$ is the number of detected instances.
@@
Multivariate Relationship Modeling
Unary Association. To establish the multivariate relationships among the detected instances in $\mathcal{Z}$, we start with the most straightforward scenario, i.e. examining the visual and textual representation of each instance (whether human or object) individually. In this unary association, its visual feature is denoted as a unary token $\mathbf{u}_{i}$, forming the set $\mathcal{U}=\{\mathbf{u}_{i}\}_{i=1}^{n}$. Meanwhile, the corresponding semantic feature $\mathbf{e}_{i}$ for each instance is obtained through the text encoder of a vision-language model (VLM). This process employs the prompt template “a photo of a/an {object}”, where {object} is the instance category $c_{i}$ detected in the first stage. Then, such category-level semantic information is combined with visual features to obtain an enhanced unary token $\mathbf{u}_{i}^{\prime}$ as,
@@
Binary Connection. Based on the unary association, we continue to investigate the binary connection between different instances. Inspired by UPT (Zhang, Campbell, and Gould 2022), a binary token set $\mathcal{G}$ = $\{\mathbf{g}_{l}\}_{l=1}^{m}$ can be constructed from the unary one $\mathcal{U}$ as follows,
@@
Following the setup in DETR (Carion et al. 2020), the structure of our binary decoder is presented in the right lower corner of Fig. 2. It includes $N$ blocks, each with cascaded self- and cross-attention layers, designed to process four input, namely content query $\mathbf{C}_{q}$, positional query $\mathbf{P}_{q}$, content key $\mathbf{C}_{k}$, and positional key $\mathbf{P}_{k}$. The attention layers follow the standard process $\operatorname{Atten}(*)$ as,
@@
Ternary Relationship. Beyond the binary connections, we further explore the ternary relationships within $\mathcal{Z}$ by incorporating the concept of object affordances into the HOI task. To determine which objects can serve as functional tools in HOIs, we query a large language model (LLM) using a prompt of “Can {X} be a tool for a human to interact with {Y}?”, where X and Y represent different object categories (e.g. bat and ball). Although the prompt does not explicitly mention affordances, the LLM exhibits a strong implicit understanding of object functionality and affordance-related reasoning. Based on LLM’s responses, we construct a set of object-tool pairs, which are stored in the knowledge bank $\mathcal{B}$ as ordered pairs of the form <<Y, X>>. To improve efficiency, these identified object-tool pairs are curated offline and used as external knowledge during training and inference.
@@
A ternary decoder, designed with a structure similar to the binary decoder but having different coefficients, is developed to leverage affordance-guided interaction from $\mathcal{T}$. Noting that each ternary token $\mathbf{t}_{o}\in\mathbb{R}^{D}$ is associated with a triplet consisting of the human token $\mathbf{u}_{i}^{\prime}$, object token $\mathbf{u}_{j}^{\prime}$ and tool token $\mathbf{u}_{k}^{\prime}$. Based on their corresponding bounding box triplet $(\mathbf{b}_{i},\mathbf{b}_{j},\mathbf{b}_{k})$, we compute the relationships between each pair of tokens, i.e. human-object, human-tool, and object-tool. These pairwise features are then concatenated and passed through an MLP to form the ternary positional matrix $\mathbf{W}\in\mathbb{R}^{r\times D}$, which serves as the positional query $\mathbf{P}_{q}$. Then, the ternary matrix $\mathbf{T}_{0}\in\mathbb{R}^{r\times D}$, concatenated by all $\mathbf{t}_{o}$, is used as the content query $\mathbf{C}_{q}$. The output of self-attention $\mathbf{A}_{i}^{\textit{t-self}}$ in $i$-th ternary decoder block is,
@@
Contextualized Prompt Learning
@@
The corresponding object labels (i.e. $c_{j}$) are used to construct a prompt in the form of “[PREFIX] person [ACT] {object}”, where [PREFIX] and [ACT] can be a sequence of learnable tokens, i.e. $\mathbf{v}]_{1}[$$\mathbf{v}]_{2}[$$\mathbf{v}]_{A}$. Each $\mathbf{v}]_{a}$ ($a\in\{1,...,A\}$) is a vector with the same dimension as word embeddings (i.e. 512 or 768 for CLIP), and $A$ is a hyperparameter specifying the number of learnable tokens. Then, they are passed through the VLM’s text encoder and a projection layer to produce the textual features $\mathbf{M}_{0}\in\mathbb{R}^{m\times C^{\prime}}$.
A new contextual decoder, whose block shares the same structure as the binary and ternary ones, is proposed to fuse regional and global contextual visual information (i.e. $\mathbf{D}$ and $\mathbf{V}_{g}$) with adaptive textual features $\mathbf{M}_{0}$ for interaction reasoning. Specifically, two decoder blocks (global and regional), i.e. $N$ = 2, are designed to sequentially inject $\mathbf{V}_{g}$ and $\mathbf{D}$ at the cross-attention layer. Meanwhile, the positional queries or keys are set to $\mathbf{0}$, since the location of instances is less relevant to the prompt. Given $\mathbf{M}_{0}$ as the input, the global visual context $\mathbf{V}_{g}$ is utilized to form the context key and value at the first cross-attention layer. It is repeated $m$ times to obtain $\mathbf{V}^{\prime}_{g}\in\mathbb{R}^{m\times C^{\prime}}$, aligned with the input’s dimensions. Thus, the output $\mathbf{M}_{1}$ of the first block can be formulated as,

### v2

# Contextualized Representation Learning for Effective Human-Object Interaction Detection

@@ Abstract
Human-Object Interaction (HOI) detection aims to simultaneously localize human-object pairs and recognize their interactions. While recent two-stage approaches have made significant progress, they still face challenges due to incomplete context modeling. In this work, we introduce a Contextualized Representation Learning that integrates both affordance-guided reasoning and contextual prompts with visual cues to better capture complex interactions. We enhance the conventional HOI detection framework by expanding it beyond simple human-object pairs to include multivariate relationships involving auxiliary entities like tools. Specifically, we explicitly model the functional role (affordance) of these auxiliary objects through triplet structures <<human, tool, object>>. This enables our model to identify tool-dependent interactions such as “filling”. Furthermore, the learnable prompt is enriched with instance categories and subsequently integrated with contextual visual features using an attention mechanism. This process aligns language with image content at both global and regional levels. These contextualized representations equip the model with enriched relational cues for more reliable reasoning over complex, context-dependent interactions. Our proposed method demonstrates superior performance on both the HICO-Det and V-COCO datasets in most scenarios. The source code is available at https://github.com/lzzhhh1019/CRL.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Human-Object Interaction (HOI) detection is a fundamental and challenging task in visual recognition that aims to model complex semantic relationships between humans and surrounding objects in a visual scene, ultimately recognizing triplets of the form <<human, action, object>>. It requires a simultaneous understanding of visual features, spatial configurations and contextual semantics to effectively identify all valid interaction patterns within given images. Therefore, other high-level semantic understanding tasks, such as activity recognition [1, 2, 3] and video comprehension [4, 5, 6] can benefit from HOI.
@@
Figure 1: Contextualized representations in HOI. (a) Tool affordances (e.g., bottle’s pourable) help distinguish complex interactions (e.g., human fill cup) from direct human–object relations (e.g., human hold cup). (b) Contextualized alignment with instance categories (e.g., cup) and corresponding visual features narrows down the potential actions to relevant ones (e.g., fill, hold).
@@
Existing HOI detection approaches can generally be divided into one-stage and two-stage methods. One-stage methods perform object detection and interaction prediction simultaneously, but adapting their decoders for various HOIs often incurs substantial training costs, sometimes requiring hundreds of GPU hours. On the other hand, two-stage methods leverage pre-trained object detectors to identify humans and objects before explicitly establishing relationships between them for interaction prediction. Thanks to its flexibility, the two-stage approach has quickly become the preferred paradigm. However, traditional two-stage HOI approaches have primarily relied on features derived from human-object pairs, such as human pose [7] or spatial information [8]. In real-world interactions, many scenarios are not solely defined by these pairs. They may also involve auxiliary tools or objects that play a crucial role in the interaction. While prior HOI studies [9] have leveraged object affordances, such modeling usually remains tied to the acted-on object and does not explicitly account for tool affordances. In contrast, we model the functional role of tools within the <<human, tool, object>> ternary. For example, as illustrated in Fig. 1(a), when a person fills a cup using a bottle, understanding the bottle’s affordance (i.e., its potential function like being pourable) offers significant insights into deciphering the nature of the interaction.
@@
Meanwhile, recent advancements in pre-trained language models like BERT [10] and GPT [11] offer rich semantic insights for understanding HOIs. Prompt Learning, as an efficient method for finetuning language models, has been extened to visual tasks by CoOp [12]. Many studies have as integrated Prompt Learning into the HOI task [13, 14, 15]. However, these approaches rely exclusively on textual data and do not integrate image information from specific samples. As a result, crucial contextual visual cues are left untapped, which are essential for accurately capturing interactions.
@@
It’s worth noting that both the affordances and the text-related appearance of a specific object act as vital forms of contextualized representation, essential for identifying complex interactions. Despite the importance, their applications in HOI has not been thoroughly explored. This gap in research greatly motivates us to delve deeper into this area. In this work, we propose a new two-stage HOI detection framework, named Contextualized Representation Learning (CRL). To be specific, unlike traditional human-object pairs, multivariate relationships (unary, binary and ternary ones) are modeled to capture tool-mediated interactions through affordance-guided triplets <<human, tool, object>>. In parallel, the contextualized representations of detected instances are injected into the process of prompt learning. Specifically, as illustrated in Fig. 1(b), we extract localized visual features from the corresponding instance regions and refine the prompt representation through a cross-attention mechanism with VLM’s global image embedding. This design allows the prompt to be conditioned on both textual semantics and instance-level visual cues, improving its capacity to capture context-dependent interactions.
@@
II Related Work
II-A HOI Detection
@@
Existing Human-Object Interaction (HOI) detection approaches can be divided into one-stage and two-stage methods. One-stage methods [16, 17, 18, 19, 20] simultaneously predict human and object boxes, categories, and interaction classes. Recent advancements in one-stage detectors, particularly those utilizing transformer architectures [21, 22, 23, 24], have shown promising performance. This end-to-end approach simplifies the inference process of HOI. However, it imposes heavy computational costs during training, which can limit its practicality.
@@
To alleviate this issue, two-stage methods [25, 26, 27, 28, 29, 30] undergone significant development recently. They typically rely on pre-trained detectors (e.g., DETR) to generate object proposals, focusing more on extracting interaction context from candidate human-object pairs. For example, UPT [8] first detects all humans and objects (i.e., unary features) and then applies self-attention to unary features and human-object pairs, effectively enhancing the confidence of positive samples. PViC [31] incorporates image features back into the representation of human-object pairs via cross-attention to compensate for missing contextual information. Nonetheless, traditional methods often treat all objects equally, failing to account for the role that tools play in interactions. In this work, we explore the concept of tool affordances to more effectively capture these relational cues.
II-B Vision-Language Models in HOI
@@
To obtain more effective HOI representations, several studies [32] explore knowledge transfer from vision-language pre-trained models (e.g., CLIP [33]). This approach not only enriches the learned representations but also improves the model’s capability in HOI recognition. GEN-VLKT [32] utilizes CLIP knowledge for interaction classification and the distillation of visual features. HOICLIP [34] uses the features obtained by the VLM visual encoder and proposes a new transfer strategy that uses visual semantic algorithms to represent action. ViPLO [16] adopts the Vision Transformer (ViT) from CLIP as its backbone and introduces a pose-conditioned graph to capture the local features of human joints. ADA-CM [35] develops a concept-guided memory mechanism to represent visual embeddings and semantic knowledge simultaneously. The success of VLMs opens up new avenues for HOI detection. CLIP4HOI[36] utilizes CLIP’s vision-language knowledge to enhance human-object interaction detection by decoupling human and object detection and adapting CLIP into a fine-grained classifier for better interaction discrimination.
II-C Prompt Learning in HOI
@@
Prompt learning has become very popular for fine-tuning VLMs on downstream tasks. Context Optimization (CoOp) [12] encodes prompt context as learnable vectors, achieving strong performance with only a few labeled samples. To better adapt to downstream tasks, Conditional Context Optimization (CoCoOp) [37] builds upon CoOp by employing a lightweight network to generate input-conditional context tokens for each image. THID [13] is the first to propose the use of learnable language prompts for the HOI detection task. CMMP [14] introduces decoupled multi-modal prompts for spatial-aware HOI detection, separating visual feature extraction and interaction classification to reduce error propagation. EZ-HOI [15] leverages LLM-generated class descriptions to guide prompt learning.
@@
Figure 2: Overall architecture of our Contextualized Representation Learning Network, consisting of Multivariate Relationship Modeling (MRM) and Contextualized Prompt Learning (CPL). MRM constructs unary, binary and ternary token sets from regional features to model HOIs. CPL builds a category-aware learnable prompt, fused with diverse contextual visual features. Their combined outputs are utilized for interaction prediction. The structure of the binary/ternary/contextual decoder is shown in the bottom right.
III Method
III-A Preliminary
Our proposed model builds upon the traditional two-stage methodology for Human-Object Interaction (HOI) detection, which first focuses on detecting instances and subsequently classifies the interactions between them. In the first stage, as shown in the upper-left part of Fig. 2, we employ an off-the-shelf object detector, such as DETR [38], to extract all instances, including humans and objects. As a result, a set of instances $\mathcal{Z}=\{z_{i}\}_{i=1}^{n}$ can be constructed with each detection $z_{i}\,=\,(\mathbf{b}_{i},s_{i},c_{i},\mathbf{u}_{i})$, consisting of the box coordinates $\mathbf{b}_{i}\in\mathbb{R}^{4}$, the confidence score $s_{i}\in[0,1]$, the detected instance $c_{i}\in\mathcal{O}$ for the category set $\mathcal{O}$, and the unary instance feature $\mathbf{u}_{i}\in\mathbb{R}^{C}$. $C$ is the embedding dimension, $n$ is the number of detected instances.
@@ III-A Preliminary
III-B Multivariate Relationship Modeling
Unary Association. To establish the multivariate relationships among the detected instances in $\mathcal{Z}$, we start with the most straightforward scenario, i.e., examining the visual and textual representation of each instance (whether human or object) individually. In this unary association, its visual feature is denoted as a unary token $\mathbf{u}_{i}$, forming the set $\mathcal{U}=\{\mathbf{u}_{i}\}_{i=1}^{n}$. Meanwhile, the corresponding semantic feature $\mathbf{e}_{i}$ for each instance is obtained through the text encoder of a vision-language model (VLM). This process employs the prompt template “a photo of a/an {object}”, where {object} is the instance category $c_{i}$ detected in the first stage. Then, such category-level semantic information is combined with visual features to obtain an enhanced unary token $\mathbf{u}_{i}^{\prime}$ as,
@@ III-B Multivariate Relationship Modeling
Binary Connection. Based on the unary association, we continue to investigate the binary connection between different instances. Inspired by UPT [8], a binary token set $\mathcal{G}$ = $\{\mathbf{g}_{l}\}_{l=1}^{m}$ can be constructed from the unary one $\mathcal{U}$ as follows,
@@
Following the setup in DETR [38], the structure of our binary decoder is presented in the right lower corner of Fig. 2. It includes $N$ blocks, each with cascaded self- and cross-attention layers, designed to process four input, namely content query $\mathbf{C}_{q}$, positional query $\mathbf{P}_{q}$, content key $\mathbf{C}_{k}$, and positional key $\mathbf{P}_{k}$. The attention layers follow the standard process $\operatorname{Atten}(*)$ as,
@@
Ternary Relationship. Beyond the binary connections, we further explore the ternary relationships within $\mathcal{Z}$ by incorporating the concept of object affordances into the HOI task. To determine which objects can serve as functional tools in HOIs, we query a large language model (LLM) using a prompt of “Can {X} be a tool for a human to interact with {Y}?”, where X and Y represent different object categories (e.g., bat and ball). Although the prompt does not explicitly mention affordances, the LLM exhibits a strong implicit understanding of object functionality and affordance-related reasoning. Based on LLM’s responses, we construct a set of object-tool pairs, which are stored in the knowledge bank $\mathcal{B}$ as ordered pairs of the form <<Y, X>>. To improve efficiency, these identified object-tool pairs are curated offline and used as external knowledge during training and inference.
@@
A ternary decoder, designed with a structure similar to the binary decoder but having different coefficients, is developed to leverage affordance-guided interaction from $\mathcal{T}$. Noting that each ternary token $\mathbf{t}_{o}\in\mathbb{R}^{D}$ is associated with a triplet consisting of the human token $\mathbf{u}_{i}^{\prime}$, object token $\mathbf{u}_{j}^{\prime}$ and tool token $\mathbf{u}_{k}^{\prime}$. Based on their corresponding bounding box triplet $(\mathbf{b}_{i},\mathbf{b}_{j},\mathbf{b}_{k})$, we compute the relationships between each pair of tokens, i.e., human-object, human-tool, and object-tool. These pairwise features are then concatenated and passed through an MLP to form the ternary positional matrix $\mathbf{W}\in\mathbb{R}^{r\times D}$, which serves as the positional query $\mathbf{P}_{q}$. Then, the ternary matrix $\mathbf{T}_{0}\in\mathbb{R}^{r\times D}$, concatenated by all $\mathbf{t}_{o}$, is used as the content query $\mathbf{C}_{q}$. The output of self-attention $\mathbf{A}_{i}^{\textit{t-self}}$ in $i$-th ternary decoder block is,
@@
III-C Contextualized Prompt Learning
@@
The corresponding object labels (i.e., $c_{j}$) are used to construct a prompt in the form of “[PREFIX] person [ACT] {object}”, where [PREFIX] and [ACT] can be a sequence of learnable tokens, i.e., $\mathbf{v}]_{1}[$$\mathbf{v}]_{2}[$$\mathbf{v}]_{A}$. Each $\mathbf{v}]_{a}$ ($a\in\{1,...,A\}$) is a vector with the same dimension as word embeddings (i.e., 512 or 768 for CLIP), and $A$ is a hyperparameter specifying the number of learnable tokens. Then, they are passed through the VLM’s text encoder and a projection layer to produce the textual features $\mathbf{M}_{0}\in\mathbb{R}^{m\times C^{\prime}}$.
A new contextual decoder, whose block shares the same structure as the binary and ternary ones, is proposed to fuse regional and global contextual visual information (i.e., $\mathbf{D}$ and $\mathbf{V}_{g}$) with adaptive textual features $\mathbf{M}_{0}$ for interaction reasoning. Specifically, two decoder blocks (global and regional), i.e., $N$ = 2, are designed to sequentially inject $\mathbf{V}_{g}$ and $\mathbf{D}$ at the cross-attention layer. Meanwhile, the positional queries or keys are set to $\mathbf{0}$, since the location of instances is less relevant to the prompt. Given $\mathbf{M}_{0}$ as the input, the global visual context $\mathbf{V}_{g}$ is utilized to form the context key and value at the first cross-attention layer. It is repeated $m$ times to obtain $\mathbf{V}^{\prime}_{g}\in\mathbb{R}^{m\times C^{\prime}}$, aligned with the input’s dimensions. Thus, the output $\mathbf{M}_{1}$ of the first block can be formulated as,